OpenStudy (anonymous):

Express the complex number in trigonometric form. -2

OpenStudy (anonymous):

I know it's in the 2nd quadrant and the angle is 63 degrees but when I do 180-63 to find it in the 2nd quadrant it gives me 117 and that's not one of my answer choices??

OpenStudy (anonymous):

@cwrw238 will you help me please?

OpenStudy (cwrw238):

let complex number be z z = -2 + 0i = r( cosA + isinA) in polar form where r = Sqrt(-2^2 + 0^1) and A = arctan 0

OpenStudy (anonymous):

2(cos 0° + i sin 0°) ?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

ok thank you for the help

OpenStudy (cwrw238):

yw